A satellite is put into orbit around the earth first by being carried on a rocket that climbes vertically to h = 217 km from the earth\'s surface and then being propelled at a velocity of v = 9.65 km/s. This initial velocity is perpendicular to the position vector from the earth\'s center to the satellite. What is the maximum distance from the earth\'s center the satellite will reach with a closed orbit around the earth?
Answer #1

i think this helps
Answer #2

According to the law of conservation of energy

\frac{1}{2}mv^{2}+mg(R+h)=\frac{1}{2}mv_{0}^{2}+mgd

but, v_{o}^{2}=\frac{GM}{d}

Then the above expression can be written as,

\frac{1}{2}mv^{2}+mg(R+h)=\frac{1}{2}(\frac{GM}{d})+mgd

2gd^{2}-d[v^{2}+2g(R+h)]+GM=0

By substituting the values in the above equation, we have

            d = 5.89 * 10^6 m